OpenStudy (anonymous):

How do you solve... 2x+3/5=x-4/2

OpenStudy (anonymous):

for x?

OpenStudy (anonymous):

x = - 13/5 13 over 5 like a fraction

OpenStudy (anonymous):

\[\frac{ 2x+3 }{ 5 }=\frac{ x-4 }{ 2}\]

OpenStudy (anonymous):

I would assume solve for x

OpenStudy (anonymous):

2x+3/5=x-4/2 2x- x = -4/2 - 3/5 x = -8/10 - 6/10 u can do the rest...

OpenStudy (anonymous):

sorry that format is heaps diff u should of posted to begin with

OpenStudy (anonymous):

just rearrange to isolate x, get like terms on one side

OpenStudy (anonymous):

sorry... didnt know how to do that format before

OpenStudy (mathstudent55):

You have a fraction equal to a fraction. Cross multiply them just as if you were solving a proportion.

OpenStudy (anonymous):

2x +3/5 =x -4/5 2x-x= -(4/5) - (3/5) x =-(26/10) x=-(13/5)

OpenStudy (mathstudent55):

2(2x + 3) = 5(x - 4) Distribute the 2 and the 4, collect like terms, and solve for x.

OpenStudy (anonymous):

So x would be 26?

OpenStudy (anonymous):

verify: 2(-13/5) + (3/5) =(-13/5) - (4/2) =-23/5

OpenStudy (mathstudent55):

2(2x + 3) = 5(x - 4) 4x + 6 = 5x - 20 26 = x Correct.

OpenStudy (anonymous):

Thank you @mathstudent55

OpenStudy (anonymous):

2x + 3 = (x - 4 / 2) * 5 2x = (x - 4 / 2) * 5 - 3 x = ((x - 4 / 2) * 5 - 3)/2 simplify here

OpenStudy (mathstudent55):

@sss1980 Read the first response by the original poster. The problem is not what the original post means.

OpenStudy (mathstudent55):

@melb94 You're welcome.